OpenStudy (anonymous):

prove: sin y+tan y/ over /1+ sec y =sin y

geerky42 (geerky42):

Like this? \(\dfrac{\sin y + \tan y}{1 + \sec y} = \sin y\)

OpenStudy (anonymous):

yes

geerky42 (geerky42):

Start with this: \[ \dfrac{\sin y + \tan y}{1+\sec y} \cdot \dfrac{1-\sec y}{1-\sec y} \] Then multiply both fractions together then simplify.

geerky42 (geerky42):

Get it?

OpenStudy (anonymous):

yep

geerky42 (geerky42):

Oh, and remember these trigonometry identify: 1−sec²x = −tan²x

geerky42 (geerky42):

Please let me know when you are stuck, ok?

OpenStudy (anonymous):

yea im bad at math so im stuck

geerky42 (geerky42):

In what step are you stuck at?

OpenStudy (anonymous):

just cant do it